### v3

@@ Abstract
Instruction tuning plays a critical role in aligning large language models (LLMs) with human preference. Despite the vast amount of open instruction datasets, naively training a LLM on all existing instructions may not be optimal and practical. To pinpoint the most beneficial datapoints, data assessment and selection methods have been proposed in the fields of natural language processing (NLP) and deep learning. However, under the context of instruction tuning, there still exists a gap in knowledge on what kind of data evaluation metrics can be employed and how they can be integrated into the selection mechanism. To bridge this gap, we present a comprehensive review on existing literature of data assessment and selection especially for instruction tuning of LLMs. We systematically categorize all applicable methods into quality-based, diversity-based, and importance-based ones where a unified, fine-grained taxonomy is structured. For each category, representative methods are elaborated to describe the landscape of relevant research. In addition, comparison between latest methods is conducted on their officially reported results to provide in-depth discussions on their limitations. Finally, we summarize the open challenges and propose the promosing avenues for future studies. All related contents are available at https://github.com/yuleiqin/fantastic-data-engineering.
Unleashing the Power of Data Tsunami: A Comprehensive Survey on Data Assessment and Selection for Instruction Tuning of Language Models
Yulei Qin1, Yuncheng Yang1,2, Pengcheng Guo1, Gang Li1, Hang Shao1, Yuchen Shi1, Zihan Xu1, Yun Gu2, Ke Li1, Xing Sun1, 1Tencent YouTu Lab, 2Shanghai Jiao Tong University Correspondence: yuleiqin@tencent.com
@@ 1 Introduction
One of the ultimate goal of developing large lnguage models (LLMs) is to unlock their potentials of generalization to unseen natural language processing (NLP) tasks. Towards this goal, a series of LLMs such as GPTs Brown et al. (2020); Achiam et al. (2023), LLaMAs Touvron et al. (2023a, b); AI@Meta (2024), and Mistrals Jiang et al. (2023a, 2024a) have delivered high-level text understanding and generation capabilities via utilizing vast amount of high-quality web and human-annotated datasets for pre-training and preference alignment Liu et al. (2023a, 2024c); Sun et al. (2024b); Edunov et al. (2019); Dong et al. (2019). During preference alignment, instruction tuning plays an important role in refining pre-trained LLMs to provide accurate, pertinent, and harmless responses on a collection of downstream tasks Wei et al. (2021); Sanh et al. (2021); Zhang et al. (2023c); Peng et al. (2023); Longpre et al. (2023); Shu et al. (2023); Jang et al. (2023); Ghosh et al. (2024); Kung and Peng (2023). For efficient and effective instruction tuning, existing studies Ouyang et al. (2022); Taori et al. (2023); Zhou et al. (2024a); Xia et al. (2024a) have noticed that improving quality of instruction tuning data (e.g., formulation of well-defined and complete contexts), rather than simply piling up instructions without analysis (e.g., exhaustive collection of open datasets), is of prioritized concerns.
Figure 1: Categorization of data assessment and selection methods for efficient LLM instruction tuning.
In this work, we aim to unify a wide array of data assessment and selection methods under the context of instruction tuning of LLMs. As revealed from the probabilistic view John and Draper (1975); Murphy (2012); Albalak et al. (2024), the statistical patterns inherent in datasets determines the modeling performance. The overall evaluation of instruction datapoints not only deciphers the distribution in various aspects (e.g., composition, task, and domain) and also help cherry-pick the most beneficial subsets for higher performance with less training cost. Through this survey, we demonstrate that: 1) existing resourceful data assessment methods can be categorized into three main perspectives: quality, diversity, and importance (see Fig. 1). 2) a systematic view of selection methods can be unified even they more or less exhibit coupling with the assessment techniques (see Fig. 2). It is noted that quality, diversity, and importance might be used interchangeably without strict discrimination in previous studies. But here we provide a rationalized organization taxonomy for structured elaboration. Despite the goal of being comprehensive, the present survey only provides details of certain typical, representative methods to avoid being tediously long. We hope the in-depth explanations and discussions on the selected methods provide insights into developing robust data assessment and selection pipelines for further studies.
@@ 1.1 Related Surveys
Liu et al. (2024d) studies the mainstream datasets for building LLMs, including the pre-training corpora, instruction tuning datasets, preference datasets, evaluation benchmarks, and traditional NLP datasets. Albalak et al. (2024) presents a systematic overview of constructing the data pipeline for language models. Any selection method, either via distribution matching or diversification, can be composed of: 1) utility function; 2) selection mechanism. During different stages of the pipeline (e.g., language filtering, data quality, domain knowledge, deduplication, toxic and explicit content removal, and data mixing), the selection method should be adjusted according to different selection objectives. Wang et al. (2024a) focuses on the data preparation for instruction tuning. Existing methods on building instruction tuning datasets include: 1) reformulating the discriminative NLP datasets into generative ones; 2) self-instruct with seed prompts; 3) prompt mapping and evol-instruct. Popular methods on dataset selection can be simply classified as: 1) system of indicators; 2) trainable LLMs; 3) powerful LLMs; and 4) small models. Guo et al. (2022) starts from the general coreset selection method in the field of deep learning and categorize all selection manners into: 1) geometry-based methods (e.g., herding, kcenter-greedy); 2) uncertainty-based methods (e.g., least confidence/entropy/margin); 3) error/loss-based methods (forgetting; GraND/EL2N; importance resampling); 4) decision boundary-based (adversarial deepfool; contrastive active learning); 5) gradient matching-based (gradient approximation towards full set); 6) bi-level optimization-based (inner loop of model optimization and outer loop of datapoint selection); 7) sub-modularity-based (e.g., graph cut; facility location); 8) proxy-based (preference of a small model on data selection). Zhou et al. (2024b) investigates the potential metrics and aspects for data quality measurement and provides a list of available tools for data evaluation. Apart from data assessment and selection methods that specifically designed for NLP or LLM applications Moore and Lewis (2010); Chen et al. (2024a); Dodge et al. (2020); Kandpal et al. (2022); Li et al. (2022); Feng et al. (2021); Lee et al. (2021); Malhotra and Bakal (2015); Liu et al. (2024e), there exist many survey studies that tackle general quality measurement in machine learning Gupta et al. (2021); Zha et al. (2023); Ehrlinger and Wöß (2022); Mohammed et al. (2024); Li et al. (2024c); Lu et al. (2023b); Dix et al. (2023); Priestley et al. (2023); Byabazaire et al. (2020); Roh et al. (2019); Sidi et al. (2012); Batini et al. (2009) for constructing safe, unbiased, and accurate datasets.
@@
Figure 2: A high-level overview of comprehensive data assessment and selection. Most analysis aspects that evaluate each datapoint and the overall dataset are categorized into three groups marked in blue italic.
@@ 1.2 Survey Scope
Although “data evaluation" has been so frequently mentioned that it appears as a cliché problem in developing machine learning algorithms, the optimal solution to establishing an overall data assessment and selection pipeline still remains an open question. Especially under the context of instruction tuning of LLMs, existing studies propose various measurement and cleaning strategies to select the “high-quality" instructions from all datapoints. However, very few studies notice that there exists no unified dimensions or aspects in measuring data “quality" where previous works tend to put emphasis on the domain-specific and task-dependent characteristics. In addition, the inherent, systematic coupling between data assessment and subset selection methods is not well demonstrated.
Under such circumstance, the present study strives to provide a comprehensive review on evaluating and decomposing massive instruction tuning datasets. We categorize the main aspects of data assessment in terms of quality, diversity, and importance. In each aspect, we provide a detailed survey on both traditional (e.g., hand-crafted indicators) and machine learning (e.g., model-based indicators) methods. Besides, the coreset sampling methods that fuses evaluation and selection are introduced separately in diversity and importance oriented subset construction. In consideration of the properties of instruction tuning, we focus on the text modality and start from classical text analysis metrics. Metrics that are either specific to instruction tuning or compatible with pre-training and preference alignment datasets are included since they all share general rules in data assessment.
@@ 2 Preliminaries
In instruction tuning, each text sample $I_{i}$ is usually composed of three parts: 1) instruction (either with or without system prompt), 2) input, and 3) response. For an off-the-shelf pre-trained LLM parameterized as $\theta$, a pre-determined instruction template is used to wrap $I_{i}$ into the prompt $p_{i}$ with special tokens like “<|im_start|>" and “<|im_end|>" for separation of roles (e.g., system, user, assistant, function, and observation) and their contents. Then, a LLM-associated tokenizer performs tokenization on the instruction prompt $p_{i}$ for a sequence of $x_{i}=\{x_{i(1)},x_{i(2)},...,x_{i(n)}\}$, where $x_{i(j)}$ denotes the $j$-th token of $x_{i}$ and $n$ is the total number of tokens. Out of simplicity, the token sequence $x_{i}$ can be simply split into two parts by the index $t$ where the content from the role assistant starts: 1) the instruction (input) part ($x_{i(<t)}$), and 2) the ground-truth response part ($x_{i(\geq t)}$).
@@
For each $x_{i}$, the model iteratively predicts the next token given $x_{i}(j)$ all previous tokens including the instruction context and the response completions up to the current token $x_{i}(<j)$.
@@
We aim at finding the most informative subset $S_{b}\subset S$ from the entire set $S$ under the given budget $|S_{b}|\leq b$. Mathematically, the selection of $S_{b}$ requires the quantitative evaluation $q(\cdot)$ on each datapoint $x_{i}$ and an elaborated sampling mechanism $\pi$:
$S_{b}^{*}=\pi(\operatorname*{arg}\operatorname*{max}_{x_{i}\in S}q(x_{i}),b),$ (2)
where $\pi(\cdot,b)$ denotes the sampling process with a maximum $b$ datapoints. With respect to the detailed implementation of $\pi$, either an iterative, greedy algorithm or a batch-wise heuristic rule can be adopted for compatibility with $q(\cdot)$. The expected benefits of such selection include: 1) the reduction of noise by ignoring those mislabeled, mismatched instruction-response pairs, 2) the re-balance of data distributions by down-sampling those easy, common, and similar examples while up-sampling hard, rare ones, and 3) the expedition of training in return for efficient iterations of LLMs.

### v4

@@ Abstract
Instruction tuning plays a critical role in aligning large language models (LLMs) with human preference. Despite the vast amount of open instruction datasets, naively training a LLM on all existing instructions may not be optimal and practical. To pinpoint the most beneficial datapoints, data assessment and selection methods have been proposed in the fields of natural language processing (NLP) and deep learning. However, under the context of instruction tuning, there still exists a gap in knowledge on what kind of data evaluation metrics can be employed and how they can be integrated into the selection mechanism. To bridge this gap, we present a comprehensive review on existing literature of data assessment and selection especially for instruction tuning of LLMs. We systematically categorize all applicable methods into quality-based, diversity-based, and importance-based ones where a unified, fine-grained taxonomy is structured. For each category, representative methods are elaborated to describe the landscape of relevant research. In addition, comparison between the latest methods is conducted on their officially reported results to provide in-depth discussions on their limitations. Finally, we summarize the open challenges and propose the promosing avenues for future studies. All related contents are available at https://github.com/yuleiqin/fantastic-data-engineering.
@@ 1 Introduction
One of the ultimate goal of developing large lnguage models (LLMs) is to unlock their potentials of generalization to unseen natural language processing (NLP) tasks. Towards this goal, a series of LLMs such as GPTs Brown et al. (2020); Achiam et al. (2023), LLaMAs Touvron et al. (2023a; b); AI@Meta (2024), and Mistrals Jiang et al. (2023a; 2024a) have delivered high-level text understanding and generation capabilities via utilizing vast amount of high-quality web and human-annotated datasets for pre-training and preference alignment Liu et al. (2023a; 2024c); Sun et al. (2024b); Edunov et al. (2019); Dong et al. (2019). During preference alignment, instruction tuning plays an important role in refining the pre-trained LLMs to provide accurate, pertinent, and harmless responses on a collection of downstream tasks Wei et al. (2021); Sanh et al. (2021); Zhang et al. (2023d); Peng et al. (2023); Longpre et al. (2023); Shu et al. (2023); Jang et al. (2023); Ghosh et al. (2024); Kung & Peng (2023). For efficient and effective instruction tuning, existing studies Ouyang et al. (2022); Taori et al. (2023); Zhou et al. (2024a); Xia et al. (2024a) have noticed that improving the quality of instruction tuning data (e.g., formulation of well-defined and complete contexts), rather than simply piling up instructions without analysis (e.g., exhaustive collection of open datasets), is of prioritized concerns.
Figure 1: Categorization of data assessment and selection methods for effective instruction tuning of LLMs.
In this work, we aim to unify a wide array of data assessment and selection methods under the context of instruction tuning of LLMs. As revealed from the probabilistic view John & Draper (1975); Murphy (2012); Albalak et al. (2024), the statistical patterns inherent in datasets determine the modeling performance. The overall evaluation of datapoints not only deciphers the distribution in various aspects (e.g., composition, task, and domain) but also helps cherry-pick the most beneficial subsets for higher performance with less training cost. Through this survey, we demonstrate that: 1) existing resourceful data assessment methods can be categorized into three main perspectives: quality, diversity, and importance (see Fig. 1). 2) a systematic view of selection methods can be unified even they more or less exhibit coupling with the assessment techniques (see Fig. 2). It is noted that quality, diversity, and importance might be used interchangeably without strict discrimination in previous studies. But here we provide a rationalized organization taxonomy for structured elaboration. Despite the goal of being comprehensive, the present survey only provides details of certain typical, representative methods to avoid being tediously long. We hope the in-depth explanations and discussions on the selected methods provide insights into developing robust data assessment and selection pipelines for future studies.
@@ 1.1 Related Surveys
Liu et al. (2024d) studied the mainstream datasets for building LLMs, including the pre-training corpora, instruction tuning datasets, preference datasets, evaluation benchmarks, and traditional NLP datasets. Their work focuses on the descriptions of dataset statistics (e.g., categorization, sources, and domains) without providing guidelines on utilization. In contrast, we emphasize the selection of instruction-tuning data for the improved downstream performance. Albalak et al. (2024) presented a systematic overview of constructing the data pipeline for language models. Any selection method, either via distribution matching or diversification, can be composed of: 1) utility function; 2) selection mechanism. During different stages of the pipeline, the selection method should be adjusted according to different selection objectives (e.g., language filtering, data quality control , domain knowledge division , deduplication, toxic and explicit content removal, and data mixing). Their work pays extra attention to the processing of the pre-training corpora while neglecting the fine-grained analysis of existing selection methods specifically designed for instruction tuning. In this case, our survey serves as an indispensable extension on the selection of instruction datasets. Wang et al. (2024a) focused on the data preparation for instruction tuning. Existing methods on building instruction tuning datasets include: 1) reformulating the discriminative NLP datasets into generative ones; 2) self-instruct with seed prompts; 3) prompt mapping and evol-instruct Popular methods on dataset selection can be simply classified as: 1) system of indicators; 2) trainable LLMs; 3) powerful LLMs; and 4) small models. Comparatively, our survey stems from the characteristics of data themselves, namely quality, diversity, and importance, for categorization of selection methods. In each category, we further provide subdivided groups by the selection philosophy, which deepens the understanding of selection for practical take-home messages.
@@
Existing surveys on general data selection also shed light on the principles of developing selection methods. Guo et al. (2022) started from the general coreset selection methods in the field of deep learning and categorize all selection manners into: 1) geometry-based methods (e.g., herding, k-center greedy); 2) uncertainty-based methods (e.g., least confidence/entropy/margin); 3) error/loss-based methods (forgetting; GraND/EL2N; importance resampling); 4) decision boundary-based (adversarial deepfool; contrastive active learning); 5) gradient matching-based (gradient approximation towards full set); 6) bi-level optimization-based (inner loop of model optimization and outer loop of datapoint selection); 7) sub-modularity-based (e.g., graph cut; facility location); 8) proxy-based (preference of a small model on data selection). Zhou et al. (2024b) investigated the potential metrics and aspects for data quality measurement. They provide a list of available tools for data evaluation. Apart from data assessment and selection methods that are specifically designed for NLP or LLM applications Moore & Lewis (2010); Chen et al. (2024a); Dodge et al. (2020); Kandpal et al. (2022); Li et al. (2022); Feng et al. (2021); Lee et al. (2021); Malhotra & Bakal (2015); Liu et al. (2024e), there exist many survey studies that tackle general quality measurement in machine learning Gupta et al. (2021); Zha et al. (2023); Ehrlinger & Wöß (2022); Mohammed et al. (2024); Li et al. (2024c); Lu et al. (2023b); Dix et al. (2023); Priestley et al. (2023); Byabazaire et al. (2020); Roh et al. (2019); Sidi et al. (2012); Batini et al. (2009) for constructing safe, unbiased, and accurate datasets. In this paper, we mainly focus on reviewing data selection methods for utilizing instruction tuning data in the context of large language models.
@@
Figure 2: A high-level overview of comprehensive data assessment and selection. The analysis aspects that apply to either individuals or the overall dataset can be categorized into three groups marked in blue italic.
@@ 1.2 Survey Scope
Although "data evaluation" has been so frequently mentioned that it appears as a cliché problem in developing machine learning algorithms, the optimal solution to establishing an overall data assessment and selection pipeline still remains an open question. Especially under the context of instruction tuning of LLMs, existing studies proposed various measurements and strategies to select the "high-quality" instructions. However, very few studies noticed that there exists no unified dimensions or aspects in measuring data "quality" where previous works tend to put emphasis on the domain-specific and task-dependent characteristics. In addition, the inherent, systematic coupling between data assessment and subset selection is not well demonstrated.
Under such circumstance, the present study strives to provide a comprehensive review on evaluating and decomposing massive instruction tuning datasets. We categorize the main aspects of data assessment in terms of quality, diversity, and importance. To reduce ambiguity, their definitions are first provided below.
Quality refers to the intrinsic value of the data. High-quality data typically satisfy two conditions: 1) The instructions are clear, accurate, and explicit in explaining the task at hand and the expected behavior of LLMs. 2) The responses are correct, coherent, and pertinent to the instructions. All the requirements and constraints specified in instructions should be met. Accordingly, most existing methods not only underline the clarity and fluency of instructions but also filter out mismatched, incorrect, and harmful responses via various judging and measuring techniques.
Diversity refers to the variety and richness of the dataset. During training, models that are exposed to datapoints under a wide range of scenarios enjoy a high level of generalization to unseen tasks. High-diversity data imply that each constituting datapoint is unique. The diverse dataset should cover different knowledge domains and task definitions with varying input contexts and output constraints. In light of this statement, existing methods that prioritize diversity all emphasize the removal of homogeneous, near-deduplicated datapoints in terms of lexical and semantic representations.
Importance refers to the impact of specific data points on the LLM’s performance. It implies the necessity of adding one datapoint into the selected subset for instruction tuning. The importance of a datapoint is up to factors such as the difficulty of following the instruction, the capability of the LLM under investigation, and the scale of the dataset to be selected. Under such circumstance, various methods of importance-based selection have been developed by measuring the difficulty of each datapoint, matching the gradients or performance of models that are trained on the selected and the full set, and investigating the dynamics of models in learning rich and scarce datasets.
In each aspect, we provide a detailed survey on both traditional (e.g., hand-crafted indicators) and machine learning (e.g., model-based indicators) methods. Besides, the coreset sampling methods that bridge evaluation and selection are introduced separately in diversity and importance oriented subset construction. In consideration of the properties of instruction tuning, we focus on the text modality and start from classical text analysis metrics. Metrics that are either specific to instruction tuning or compatible with pre-training and preference alignment are included since they all share general rules in data assessment.
@@
1.3 What’s Beyond the Survey Scope
The instruction tuning methods for de-bias and fairness of LLMs are not covered in the present study. We acknowledge that the bias and fairness aspects of instruction data are valued in developing responsible LLMs Gallegos et al. (2024); Chu et al. (2024); Li et al. (2023d). However, they are beyond the scope of the present study for the following reasons:
  • Most data selection methods on instruction tuning do not even notice the bias or fairness of data. They are in lack of explicitly designing steps to reduce negative impacts of biased data.
  • Existing fairness tuning methods follow their own definitions of quality and diversity in terms of gender, race, religion, profession, age, and political ideology. It is difficult to bring in all their corresponding evaluation techniques under our selection taxonomy.
  • The bias and fairness not only intersect with diversity but also with quality control, which is not an inclusive concept. For example, to control the harmful contents like hatred and racism responses, reward models or GPT scoring from quality-based selection can be used because the publicly released reward models and LLMs are already aligned with human preference.
  • The evaluation of bias and fairness is not covered in existing studies on data selection, which makes it difficult to validate the effectiveness of selection techniques in improving the fairness of LLMs.
However, we emphasize that bias and fairness should be specifically handled because instruction-tuned LLMs tend to exhibit more bias than the pre-trained LLMs Itzhak et al. (2024). Therefore, we believe the selection for a less-biased dataset would be a promising future direction in developing comprehensive selection methods.
@@ 2 Preliminaries
In instruction tuning, each text sample $I_{i}$ is usually composed of three parts: 1) instruction (either with or without system prompt), 2) input (can be empty), and 3) response.
For an off-the-shelf pre-trained LLM parameterized as $\theta$, a pre-determined instruction template is used to wrap the $I_{i}$ into a textual prompt $p_{i}$ with special tokens like "<|im_start|>" and "<|im_end|>" for separation of roles (e.g., system, user, assistant, function, and observation) and their contents. It is noted that the prompt template for organizing a text sample $I_{i}$ differs across model families. Especially, the special tokens are unique to the tokenizers and play an important role in differentiating multi-turn conversations and stopping model generations. Then, a LLM-associated tokenizer performs tokenization on the textual prompt $p_{i}$ for a sequence of token-id integers: $x_{i}=[x_{i(1)},x_{i(2)},...,x_{i(n)}]$, where $x_{i(j)}$ denotes the $j$-th token of $x_{i}$ and $n$ is the total number of tokens. Out of simplicity, the token sequence $x_{i}$ can be simply split into two parts by the index $t$: 1) the instruction part ($x_{i(<t)}$) that is fed into the LLM without being involved in loss computation, and 2) the ground-truth response part ($x_{i(\geq t)}$) that is expected as the LLM output via minimizing its language modeling loss. Without losing generality, we use the terminology of "datapoints" throughout the paper to refer to the tokenized samples $x_{i}$.
The process of processing an instruction sample $x_{i}$ is illustrated in Fig. 3. We choose one example from the Alpaca dataset Taori et al. (2023) and perform Qwen tokenization Yang et al. (2024) for demonstration.
Figure 3: The pre-processing of an instruction dataset includes: 1) template wrapping, and 2) tokenization. In the first step, we wrap the raw texts $I_{i}$ with a pre-defined chat template into the textual prompts $p_{i}$. In the second step, we perform tokenization on $p_{i}$ with the LLM-associated tokenizer for the datapoint $x_{i}$. Given the index $t$ for indicating where the loss mask of language modeling starts taking effect, we split $x_{i}$ into $x_{i(<t)}$ and $x_{i(\geq t)}$, respectively denoting the instruction part (input) and the response part (output).
@@
For each $x_{i}$, given all previous tokens $x_{i(<j)}$ , the model iteratively predicts the next token $x_{i(j)}$ at the $j$-th index. Note that the previous tokens here include both the instruction context part $x_{i(<t)}$ and the response completions up to the current token $x_{i(\geq t,<j)}$
@@
We aim at finding the most informative subset $S_{b}\subset S$ from the entire set $S$ under the given budget $|S_{b}|\leq b$. Mathematically, the selection of $S_{b}$ requires: 1) a quantitative evaluation function $q(\cdot)$ that assesses each datapoint $x_{i}$, and 2) an elaborated sampling mechanism $\pi$ that determines the rules of selection:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{S_{b}=\pi(S,b,q).}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ (2)
With respect to the detailed implementation of $\pi$, either an iterative, greedy algorithm or a batch-wise heuristic rule can be adopted for compatibility with $q(\cdot)$. For example, the greedy sampling is represented as:
$S_{b}=\pi_{\text{greedy}}(S,b,q)=\operatorname*{arg}\operatorname*{max}_{S^{\prime}\subseteq S,|S^{\prime}|\leq b}\sum_{x_{i}\in S^{\prime}}q(x_{i}).$ (3)
The greedy algorithm $\pi_{\text{greedy}}$ consistently selects datapoint $x_{i}$ with the highest score $q(x_{i})$ until the budget target $b$ is met. The heuristic sampling algorithm, on the other hand, chooses each datapoint $x_{i}$ according to a pre-defined rule. For example, the probability derived by the score $q(x_{i})$ can be employed as reference:
$S_{b}=\pi_{\text{prob}}(S,b,q)$ $=\text{Sample}(S,b,p),$ (4)
$p(x_{i})$ $=\frac{q(x_{i})}{\sum_{x_{j}\in S}q(x_{j})},$
where $\text{Sample}(S,b,p)$ performs sampling $b$ times upon the normalized probability distribution over $q(\cdot)$. More advanced sampling techniques $\pi$ can be developed in accordance with the domains and tasks at hand.
Such data assessment and selection paradigm is expected to bring about the following benefits: 1) the reduction of noise by ignoring those mislabeled, mismatched instruction-response pairs, 2) the re-balance of data distributions by down-sampling those easy, common, and similar examples while up-sampling hard, rare, and unique ones, and 3) the expedition of training in return for efficient optimization of LLMs.